OpenStudy (anonymous):

Choose the correct simplification of 5^4/5^2 1/25 10 25 1/10

OpenStudy (anonymous):

5^4 / 5^2 Dividing exponents means subtract the bottom exponent from the top one if both modify the same number. So since 5 = 5 then: 5^4 / 5^2 = 5^2 = 25.

OpenStudy (anonymous):

5^ (4-2) = 5^(2) = 25

OpenStudy (anonymous):

Take note of the rules of exponents: a^m/a^n = a^(m-n).

OpenStudy (anonymous):

Thank you!!